OpenStudy (anonymous):

Help! How do I solve this linear equation? 6x-y+z=0 4x+2y-3z=18 x-3y+2z=-15

OpenStudy (unklerhaukus):

can you use matrices?

OpenStudy (anonymous):

i can. im teaching myself.

OpenStudy (unklerhaukus):

\[\left(\begin{array}\\6&-1&1\\4&2&-3\\1&-3&2\end{array}\right)\left(\begin{array}\\x\\y\\z\end{array}\right)=\left(\begin{array}\\0\\18\\-15\end{array}\right)\]

OpenStudy (anonymous):

right...now what do i do?

OpenStudy (unklerhaukus):

do you know the valid row operations you can apply?

OpenStudy (unklerhaukus):

you can multiply any row by a constant \[R_n\rightarrow \alpha R_n\]simply multiply each of the 5 terms in the top row by a scalar

OpenStudy (anonymous):

You are really confusing me

OpenStudy (unklerhaukus):

you can swap an entire row with another \[R_n\longleftrightarrow R_m\]

OpenStudy (unklerhaukus):

maybe matrices isn't a good idea then

OpenStudy (unklerhaukus):

the other operation you can do is add or subtract one row from another\[R_n\rightarrow R_n\pm R_m\]

hero (hero):

I know matrices and I know for a fact that @UnkleRhaukus is confusing you

hero (hero):

You can't teach someone something for the first time that way

OpenStudy (unklerhaukus):

what is the easy way to solve this

hero (hero):

I'm saying, you can use matrices, but your approach to demonstrating it will most likely confuse

hero (hero):

I'm going to avoid saying "teach" here

OpenStudy (unklerhaukus):

add all the equations together what do you get (6x-y+z)+(4x+2y-3z)+(x-3y+2z)=0+18-15

OpenStudy (unklerhaukus):

\[(6x+4x+x)+(-y+2y-3y)+(z-3z+2z)=3\]

OpenStudy (unklerhaukus):

ok that dosent work either, i still cant see an simply way to do this

hero (hero):

Drawing button ?

OpenStudy (anonymous):

ive figured it out

OpenStudy (unklerhaukus):

did you get x=1

hero (hero):

How?

OpenStudy (anonymous):

i added

OpenStudy (anonymous):

x=1 y=4 and z=-2

hero (hero):

You can't just "add".

OpenStudy (unklerhaukus):

thats right, i want to know your steps helpmewithmath2012

hero (hero):

You have to perform a sequence of operations

hero (hero):

I don't think she/he will show them

OpenStudy (unklerhaukus):

just outline the method i want to know which rows to add

OpenStudy (unklerhaukus):

i want to know your method too hero

OpenStudy (anonymous):

i solved y by getting 11x+2y=3 y=4 but idk how i got x and z

OpenStudy (unklerhaukus):

did you do it in you head or use paper @helpmewithmath2012 ?

OpenStudy (anonymous):

my head so now im trying to figure out what i did

OpenStudy (turingtest):

do you know how to find the determinant of a matrix?

OpenStudy (zepp):

Can't you simply use row reduction until you get something like? @UnkleRhaukus